OpenStudy (anonymous):

If a and b (not equal zero) are the roots of the equation x^2+ax+b=0, then find least value of x^2+ax+b (x belongs R)..?

OpenStudy (anonymous):

sum of roots=a+b=-a hence b=-2a.........(1) product of roots=ab=b canceling b b/s a=1 put a=1 in equation (1) b=-2(1) b=-2

OpenStudy (anonymous):

The quadratic formula says the roots are \[x_1,x_2=\frac{-a\pm\sqrt{a^2-4b}}{2}\] or \[\large\begin{cases} a=\dfrac{-a+\sqrt{a^2-4b}}{2}\\\\ b=\dfrac{-a-\sqrt{a^2-4b}}{2} \end{cases}\] The first equation gives \[3a=\sqrt{a^2-4b}~~\iff~~9a^2=a^2-4b~~\iff~~a^2=-\frac{b}{2}\] which means \(b<0\), otherwise \(a\) is complex. Substituting into the second equation in the system, we can have \[\large -2a^2=\frac{-a-\sqrt{a^2-4(-2a^2)}}{2}\] We should be able to solve for \(a\) now. \[\begin{align*}-4a^2&=-a-\sqrt{a^2+8a^2}\\ 4a^2&=a+3\sqrt{a^2} \end{align*}\] Assume \(a>0\) for now, then \(\sqrt{a^2}=|a|=a\), so \[4a^2=4a\] which has roots \(a=0\) and \(a=1\). If we assumed \(a<0\), we'd have \[4a^2=-2a\] which has complex solutions for \(a\) (which presumably aren't desirable). The roots are said to be non-zero, so \(a=1\), which makes \(b=-2\).

OpenStudy (anonymous):

Oops, \(4a^2+2a\) doesn't have complex roots, just negative ones. My mistake

OpenStudy (anonymous):

hence to find the value of x from x^2+x(1)-2=0 middle term splitting gives x^2+2x-x-2=0 x(x+2)-(x+2)=0 (x-1)(x+2)=0 x-1=0 & x+2=0 x=+1,-2 hence least value of x=-2

OpenStudy (anonymous):

If a and b are roots, then when you substitute them in, you get an expression that must be true. So, a^2 + a^2 + b = 0 b = -2a^2 and b^2 + ab + b = 0----->b=0 or b + a + 1 = 0 So, we can substitute the first expression into the second and get -2a^2 + a + 1 = 0. Multiply by -1 and get 2a^2 - a - 1 = 0 (2a + 1)(a-1) = 0 a=-1/2 or a=1 so a = -1/2 is the least value. Substitute back into first expression to get b=-2(-.5)^2 or b = -1/2

OpenStudy (anonymous):

I'm not sure what you mean by the least value of x^2+ax+b?? Do you mean find the minimum value of f(x) = x^2 - 1/2x - 1/2, substituting a and b back in? If so, the minimum value is on the axis of symmetry for this parabola. Axis of symmetry formula is x=-b/2a or x= (.5)/2 which is x=1/4. substitute that in and get (1/4)^2 - 1/2(1/4) -1/2 which is -9/16. So minimum point has coordinates (1/4 , -9/16 ) Hope this helps..!!

OpenStudy (anonymous):

Since a,b are roots of x^2+ax+b (x-a)(x-b)=x^2+ax+b x^2-(a+b)x+ab=x^2+ax+b => a=-(a+b) => b=-2a...i and =>ab=b => b(a-1)=0...ii from i and ii, (-2a)(a-1)=0 either -2a=0 => a=0 or a-1= 0 =>a=1 Now, when a=0 then b=0 and when a=1 then b=-2 SO, two possible solution of a and b are a=0,b=0 and a=1,b=-2

OpenStudy (anonymous):

case a=0,b=0 x^2+ax+b=x^2 since x^2>=0 , its least value is 0. case a=1 and b=-2 x^2+ax+b=x^2+x-2=x^2+2*x*1/2+(1/2)^2-(1/2)^2-2 = (x+1/2)^2 - 9/4 since (x+1/2)^2-9/4 >=-9/4 , its least value is -9/4 Thus, the least value of the given function is -9/4

OpenStudy (anonymous):

can someone explain how this is possible that (x+1/2)^2-9/4 >=-9/4 , its least value is -9/4

OpenStudy (anonymous):

Since square of a number is always greater than or equal to 0. (x+1/2)^2 >= 0 (x+1/2)^2-9/4 >=-9/4 hence its least value is -9/4

OpenStudy (anonymous):

to aryandecoolest if a=-1/2 &b=-1/2,then it should satisfy the values of sum of roots & product of roots hence a+b= -a & ab= b but (-1/2)-(1/2)=-1\[\neq1/2\] =-a & ab=(-1/2)*(-1/2)=(1/4)\[\neq-1/2=b\] hence a=-1/2 &b=-1/2 must be rejected........

Parth (parthkohli):

The least value of a quadratic expression is given by\[\dfrac{-D}{4a}\]Where \(D\) is the discriminant. We can also find the values of \(a,b\) by using sum and product formulas.

OpenStudy (anonymous):

i am not able to understand all the people are approaching this problem in different ways and giving different solutions for the problem.... which is the correct way........

Parth (parthkohli):

Correct: all of them. Easiest: subjective. Just pick any one of them and enjoy.

OpenStudy (anonymous):

if a=1 &b=-2 then putting the value in \[x ^{2}+ax+b=0\] we get, \[x ^{2}+x(1)-2=0\] & doing middle term splitting we get \[x^{2}+2x-x-2=0\] x(x+2)-(x+2)=0 (x-1)(x+2)=0 x=1&x=2 and \[+2\neq-2\] which is the root b=-2 why??? somebody please explain....

OpenStudy (anonymous):

very dramatic!!!

OpenStudy (anonymous):

funny!!!

Parth (parthkohli):

You have to put \(a\) for \(x\), not \(a\)

OpenStudy (anonymous):

ok i got it........